A sample of ten people from an outer Sydney suburb travelled the following distances each day:-     
11.4    16.6      35.4   22.0    30.5    45.9    23.8     28.6     41.0       26.6
Find the difference between the median and the mean of the sample (correct to 2 dec pl)?
Use {median - mean} as the order of the difference.
Answer #1

Mean=sum of all the observations/No: of observations

Sum of all the observations=11.4+16.6+35.4+22+30.5+45.9+23.8+28.6+41+26.6. =281.8. (282)

No: of observations= 10

Mean. =282/10

=28.2

Median. = (N/2)th item

Arranging observations in ascending order:

11.4;16.6;22;23.8;26.6;28.6;30.9;35.4;41;45.9

Median=N/2th item=10/2th item

=5th item

=26.6(27)

Difference between median and mean=27-28.2

=-1.2